OpenStudy (anonymous):

Enter the expression 2cos2(θ)−1, where θ is the lowercase Greek letter theta.

OpenStudy (anonymous):

\[cos(2\theta)\]

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

It's my pleasure.

OpenStudy (anonymous):

How would you get it though?

OpenStudy (anonymous):

Sorry for the late reply.

OpenStudy (anonymous):

\[2\cos^{2}\theta-1=2\cos^{2}\theta-(\sin^{2}\theta+\cos^{2}\theta)=\cos^{2}\theta-\sin^{2}\theta\] \[=cos\theta cos\theta-sin\theta sin\theta=cos(\theta+\theta)=cos(2\theta)\]